OpenStudy (anonymous):

If sin A = 2/3 where 0° < A < 90°, what is the value of sin 2A

OpenStudy (anonymous):

Looking at this question makes my brain hurt :(

OpenStudy (anonymous):

if sin 2A = 2 * sin a * cos a so find cos with cos(square) + sin(square) = 1

OpenStudy (anonymous):

that would be x(square) + (2/3)(square) = 1

OpenStudy (anonymous):

that gives us cos a = squareroot(5)/3

OpenStudy (anonymous):

now plug it in. sin 2A = 2 * (2/3) * (squareroot(5)/3) sin 2A = approx. 0.994

OpenStudy (anonymous):

:P just took this a while ago. hope it's right. good luck